# Diffusion Models for Image Restoration and Enhancement: A Comprehensive Survey

### v2

@@ 3 Diffusion model-based Image Restoration Methods
According to whether the diffusion models (DMs) are training-free for IR, we can preliminarily classify the DM-based IR methods into two categories, i.e., supervised DM-based methods (Saharia et al., 2022b; Whang et al., 2022; Jin et al., 2022; Qiu et al., 2023; Wang et al., 2023c; Zhou et al., 2023; Jiang et al., 2023b; Chen et al., 2023d; Chan et al., 2023; Yi et al., 2023; Gou et al., 2023; Xu et al., 2023; Wang et al., 2023h, 2024d; Zhang et al., 2023c), and zero-shot DM-based methods (Choi et al., 2021; Kawar et al., 2022a; Chung et al., 2023b; Fei et al., 2023; Song et al., 2023; Zhu et al., 2023; Mardani et al., 2023; Fabian et al., 2023; Song et al., 2021). Particularly, the supervised DM-based IR methods entail training the diffusion model from scratch with paired distorted/clean images of IR datasets. Unlike previous GAN-based methods (Yuan et al., 2018; Dinh et al., 2022; Wang et al., 2020b) that directly take distorted images as input, DM-based IR employ the well-designed conditional mechanism to incorporate the distorted images as guidance during the reverse process. Despite its promising texture generation results, this approach encounters two notable limitations: 1) training the diffusion model from scratch relies on a large quantity of paired training data; 2) collecting paired distorted/clean images in the real world is challenging. In contrast, zero-shot DM-based methods offer an appealing alternative, requiring only distorted images and dispensing with the need for retraining diffusion models. Instead of acquiring the restoration capability from the training datasets of IR, it excavates and exploits the structure and texture priors from the pre-trained diffusion models for image restoration. The core idea stems from the intuition that the pre-trained generative models can be viewed as the structure and texture repository, constructed using real-world datasets with large amounts of data, such as ImageNet (Russakovsky et al., 2015b) and FFHQ (Karras et al., 2019). Consequently, an essential challenge faced by zero-shot DM-based IR methods is: how to extract the corresponding perceptual priors while preserving the data structure from distorted images. In the subsequent subsections, we first briefly review the representative supervised DM-based IR method: SR3 (Saharia et al., 2022b), and zero-shot DM-based IR method: ILVR (Choi et al., 2021). Then we further classify these two types of methods from the perspectives of conditional strategy, diffusion modeling, and framework, which are summarized in Table 1 and Table 2 of the Supplementary, respectively. In addition, the overall taxonomy of diffusion models is illustrated in Fig. 2
@@ 3.1 SR3 – Representative Supervised DM for IR
As depicted in Fig. 3, SR3 follows the typical DDPM (Ho et al., 2020) framework and utilizes the U-Net model as the noise predictor. Given the low-resolution (LR) image $y$, SR3 initially up-samples it to the desired resolution using bicubic interpolation. Subsequently, it concatenates the super-resolved LR image $y$ with the denoised output $x_{t}$ at $t$ time step, serving as the input of the diffusion model to predict the noise for the t1t-1 step. When reaching $t=0$, the diffusion model can deliver an upsampled high-quality image $x_{0}$ of $y$ as $x_{0}\approx x$.
@@ 3.2 ILVR – Representative Zero-shot DM for IR
Although the supervised DM-based IR methods have exhibited remarkable performance, the training process entails substantial computational cost and large-scale paired datasets, which might be prohibitive for some researchers. To circumvent this, zero-shot DM-based IR (Choi et al., 2021; Lugmayr et al., 2022; Kawar et al., 2021; Chung et al., 2022b; Zhang et al., 2022c) is proposed to exploit the intrinsic knowledge within the pre-trained diffusion models. Concretely, it is observed that pre-trained diffusion models for image generation, trained with a large number of natural images, encapsulate considerable prior knowledge about rich textures. And thus, these pre-trained diffusion models can be considered as repositories of texture information. The exploration of reusing such prior knowledge for training-free image restoration is an emerging and promising direction in the low-level vision field.
As the initial work, Choi et al.(Choi et al., 2021) introduces the iterative latent variable refinement (i.e., ILVR) method, which leverages the unconditional diffusion model to enable the training-free conditional generation for image SR and image translation. The pivotal innovation of ILVR involves substituting the low-frequency components in denoised output with their counterpart from the reference image. This substitution process, illustrated in Fig. 4, ensures structure and semantic consistency between the generated and reference images, thereby facilitating the conditional generation. Particularly, given a reference image $y$ (e.g., the distorted image in IR), at time step $t$, the ILVR predicts the denoised result for time t1t-1 using the following formula:
@@
Here, $\Phi_{N}$ denotes the low-pass filter designed to bypass the low-frequency component from input, and $y_{t-1}\sim q(y_{t-1}|y)$ is the diffused state of $y$ at t1t-1 steps. Following ILVR, most zero-shot DM-based IR methods (Chung et al., 2023b; Song et al., 2023; Kawar et al., 2022a; Wang et al., 2023g; Fabian et al., 2023; Feng et al., 2023; Zhang et al., 2023b) predominantly focus on enhancing the refinement strategies within the sampling process, thereby training-free.
@@
Motivated by SR3 (Saharia et al., 2022b), numerous studies have endeavored to optimize the supervised DM-based IR framework, focusing on enhancing the condition strategy and exploring potential and more efficient generation spaces. With respect to the condition strategy, we categorize these studies into three types based on the conditions: 1) low-quality reference image, 2) pre-processed references, and 3) revising diffusion process. Regarding generation space, the supervised DM-based IR methods can be classified into three distinct groups: image space, residual space, and latent space. If not mentioned, the majority of studies generate the restored images within image space, where the structures and textures are required to be generated directly. In contrast, the residual-space diffusion model focuses on reconstructing the residuals between the low-quality image and its corresponding high-quality image, which simplifies the complexity of generating the whole image. Latent-based methods utilize a well-designed encoder to transform the image into a compact latent space for generation, thereby improving the generation efficiency. This section will elucidate existing studies on supervised DM-based IR in terms of the above three conditional strategies and the last two generation spaces.
@@ 3.3.1 Condition with Low-quality Reference Image
As highlighted in Sec. 3.1, incorporating distorted images as conditions is indispensable and crucial for supervised DM-based IR (Chen et al., 2023c; Wang et al., 2023d; Li et al., 2024b). The SR3 method has shown that substantial performance can be achieved through a simple concatenation operation. It utilizes a direct concatenation of low-quality reference images with the denoised result at the t1t-1 step as the condition for noise prediction at the $t$ step. With the same condition strategy, Saharial et al.(Saharia et al., 2022a) propose a unified diffusion model, termed Palette, for image-to-image translation tasks, which achieves excellent performance on image colorization, inpainting, uncropping, and JPEG artifact removal. Furthermore, they investigate the effects of different optimization objectives for sample diversity and highlight the pivotal role of self-attention within U-Net for the diffusion model. Despite their effectiveness, the aforementioned methods are constrained since they only support a fixed resolution for IR once trained. To adapt the diffusion model for the real-world IR with an arbitrary size, Özdenizci et al.(Özdenizci and Legenstein, 2023) divide the degraded image and corresponding sampling results $x_{t}$ into several overlapped patches, and then utilize the patch-wise concatenation as the inputs of the diffusion model for noise prediction. Additionally, to address the inconsistency problem caused by different sampling patches in the overlapped region, this work introduces the mean estimated noise for each pixel within the overlapped region. To enhance the quality of generated images, Ho et al.(Ho et al., 2022) introduce the three-layer cascaded diffusion model based on the SR3 backbone. The first diffusion model is exploited to achieve the class-conditioned low-resolution image generation, and two additional diffusion models are cascaded to super-resolve the low-resolution generated image, resulting in a higher-resolution and more realistic generated image. In addition to the DDPM-based approaches, there is another work (Dos Santos et al., 2022) that explores the various variants of the continuous diffusion model SDE (Song et al., 2020b) for the face super-resolution using predictor-corrector sampling.
@@ 3.3.2 Condition with Pre-processed References
Even though direct concatenation with the low-quality image shows promising performance, the artifacts in low-quality images inevitably cause harmful effects on the generation of the diffusion model, especially for severe and diverse distortions. To mitigate this issue, several studies (Yin et al., 2023; Wang et al., 2023e; Shang et al., 2023a; Niu et al., 2023; Huang et al., 2017; Moser et al., 2023a; Guo et al., 2023b; Wang et al., 2024b; Wu et al., 2023) strive to enhance the condition by preprocessing low-quality images with either jointly trained modules or pre-trained restoration networks. As depicted in Fig. 5, these works can be grouped into two categories based on the preprocessing strategies, i.e., condition with pre-processed reference image and feature.
@@
Pre-processed reference image. To alleviate the side effects of artifacts in low-quality images, CDPMSR (Niu et al., 2023) exploits existing super-resolution models, e.g., RCAN (Zhang et al., 2018b), SwinIR (Liu et al., 2021), EDSR (Lim et al., 2017b), to enhance the low-quality image, thereby providing a high-quality and more reliable condition for diffusion model. Additionally, it eschews the stochastic sampling during the reverse process in favor of a deterministic denoising process, yielding superior image quality and faster inference. Notably, the pre-processing reference image can serve as not only an enhanced condition but also an initially well-restored image. Therefore, ResDiff (Shang et al., 2023a) utilizes a pre-trained CNN to generate a low-frequency content-abundant image as the initially restored image and exploits the conditional diffusion model to further generate the residuals between the pre-processed distorted image and its corresponding clean image. In contrast, DehazeDDPM (Yu et al., 2023a) tries to add a physics-aware image as an auxiliary condition to achieve complex dehazing task. Thus they generate the transmission map, and haze-free image $I$ through physical modeling. Then, these preprocessed images are then incorporated into diffusion through dynamic fusion.
@@
Pre-processed reference feature. Another popular line is to use the feature of the reference image as the condition for the diffusion model (Varanka et al., 2024; Guo et al., 2023a; Gao et al., 2023b; Jin et al., 2024; Xiao et al., 2023; Wang et al., 2023e; Liu et al., 2024d). IDM (Gao et al., 2023b), striving for continuous image super-resolution, first extracts the initial features of the low-resolution image with EDSR (Lim et al., 2017b). Then, the initial features are downsampled to multiple scales, which are utilized as the conditions for different upsampling layers in the diffusion model, with the intent of refining the implicit representation. In contrast, ShadowDiffusion (Guo et al., 2023a) leverages a pre-trained transformer backbone to extract the degradation prior (i.e., the degradation-related features) from the distorted reference image. This extracted degradation prior is exploited as the auxiliary to refine the generated shadow mask and serves as the condition for shadow-free image generation.
@@
Figure 6: The architecture of IR-SDE (Luo et al., 2023b), where the forward diffusion process approximately models the image degradation processes. The final output of the forward process is equal to a low-quality image with random noise.
@@ 3.3.3 Condition by Revising Diffusion Process
Notably, the above-mentioned supervised DM-based IR methods introduce the conditions by modifying the network while preserving the diffusion process from DDPM (Ho et al., 2020). However, this requires the generation process to start from the noise. To obviate this, some studies (Luo et al., 2023b, c, 2024; Yang et al., 2023a; Zhao et al., 2023b; Wang et al., 2024c, a; Cao et al., 2022; Delbracio and Milanfar, 2023; Xiao et al., 2023; Mirza et al., 2023) condition the diffusion model by modifying the diffusion process, such that the diffused output $x_{T}$ (i.e., the start point during the reverse process) approximates the low-quality image corrupted with few Gaussian noises. As shown in Fig. 6, Luo et al.(Luo et al., 2023b) modify the forward process with mean-reverting SDE to model the unified degradation process of IR as:
@@
where $\theta_{t}$ and $\sigma_{t}$ are the time-dependent parameters. $\mu$ denotes the distorted image and $x$ represents its corresponding high-quality image. With the mean-reverting SDE, this work successfully models the image degradation and restoration processes with the modified forward and reverse processes, respectively. This avoids the generation from pure noise and achieves better resto-ration performance. Building on IR-SDE (Luo et al., 2023b), the same team introduces the Refusion (Luo et al., 2023c), which further refines IR-SDE (Luo et al., 2023b) by optimizing aspects such as network architecture, noise level, and denoising steps, etc. To reduce the computation cost, Refusion introduces the U-Net compression strategy, thereby enabling efficient sampling in the latent space.
With the same purpose, Xie et al (Xie et al., 2023) redefine the diffusion process such that the sampling starts from the noisy image. Taking into account the diversity of noise, they derive three separate diffusion processes for the removal of Gaussian, Gamma, and Poisson noises, respectively. In contrast, InDI (Delbracio and Milanfar, 2023) introduces the continuous diffusion process as:
@@
where $x$ and $y$ are the high-quality images and their corresponding low-quality counterparts. $\epsilon$ is a small constant. It can be interpreted as the step-wise interpolation of high-quality and low-quality images at time step $t$, which decomposes the original single-step prediction of supervised image restoration into several small steps, effectively circumventing the regression-to-the-mean effects often found in conventional supervised image restoration. Different from the diffusion process in the spatial domain, HFS-SDE (Cao et al., 2022) reformulate the diffusion process for magnetic resonance (MR) reconstruction at frequency space. In this approach, the forward process progressively adds noise into high-frequency space, resulting in the final $x_{T}$ that is composed of high-frequency noise and low-frequency data. During the reverse process, HFS-SDE employs Predictor-Corrector (PC) method (Song et al., 2020b) for sampling.
Figure 7: The architecture of Predict and Refine model (Whang et al., 2022). The predictor $g_{\theta}$ generates the initial prediction of a clean image, while residual information is modeled by diffusion process.

### v3

@@ 3 Diffusion model-based Image Restoration Methods
According to whether the diffusion models (DMs) are training-free for IR, we can preliminarily classify the DM-based IR methods into two categories, i.e., supervised DM-based methods (SR3, ; deblur-DPM, ; jin2022shadowdiffusion-driven, ; qiu2023diffbfr, ; wang2023exploiting-stable-sr, ; zhou2023pyramid-pydiff, ; jiang2023low-WCDM, ; chen2023hierarchical-HIdiff, ; chan2023sud, ; yi2023diff-retinex, ; gou2023exploiting-zeroair, ; xu2023stage, ; wang2023exposurediffusion, ; wang2024enhancing, ; zhang2023ultrasound-drus, ), and zero-shot DM-based methods (choi2021ilvr, ; DDRMDRM, ; dps, ; GDP, ; iigdm, ; zhu2023denoising-diffpir, ; mardani2023variational-red-diff, ; fabian2023diracdiffusion, ; SIM-SGM, ). Particularly, the supervised DM-based IR methods entail training the diffusion model from scratch with paired distorted/clean images of IR datasets. Unlike previous GAN-based methods (yuan2018unsupervised-ganr1, ; dinh2022hyperinverter-ganr6, ; wang2020transformation-ganr2, ) that directly take distorted images as input, DM-based IR employ the well-designed conditional mechanism to incorporate the distorted images as guidance during the reverse process. Despite its promising texture generation results, this approach encounters two notable limitations: 1) training the diffusion model from scratch relies on a large quantity of paired training data; 2) collecting paired distorted/clean images in the real world is challenging. In contrast, zero-shot DM-based methods offer an appealing alternative, requiring only distorted images and dispensing with the need for retraining diffusion models. Instead of acquiring the restoration capability from the training datasets of IR, it excavates and exploits the structure and texture priors from the pre-trained diffusion models for image restoration. The core idea stems from the intuition that the pre-trained generative models can be viewed as the structure and texture repository, constructed using real-world datasets with large amounts of data, such as ImageNet (russakovsky2015imagenet, ) and FFHQ (FFHQ, ). Consequently, an essential challenge faced by zero-shot DM-based IR methods is: how to extract the corresponding perceptual priors while preserving the data structure from distorted images. In the subsequent subsections, we first briefly review the representative supervised DM-based IR method: SR3 (SR3, ), and zero-shot DM-based IR method: ILVR (choi2021ilvr, ). Then we further classify these two types of methods from the perspectives of conditional strategy, diffusion modeling, and framework, which are summarized in Table 1 and Table 2 of the Supplementary, respectively. In addition, the overall taxonomy of diffusion models is illustrated in Fig. 2
@@ 3.1 SR3 – Representative Supervised DM for IR
As depicted in Fig. 3, SR3 follows the typical DDPM (DDPM, ) framework and utilizes the U-Net model as the noise predictor. Given the low-resolution (LR) image $y$, SR3 initially up-samples it to the desired resolution using bicubic interpolation. Subsequently, it concatenates the super-resolved LR image $y$ with the denoised output $x_{t}$ at $t$ time step, serving as the input of the diffusion model to predict the noise for the t1t-1 step. When reaching $t=0$, the diffusion model can deliver an upsampled high-quality image $x_{0}$ of $y$ as $x_{0}\approx x$.
@@ 3.2 ILVR – Representative Zero-shot DM for IR
Although the supervised DM-based IR methods have exhibited remarkable performance, the training process entails substantial computational cost and large-scale paired datasets, which might be prohibitive for some researchers. To circumvent this, zero-shot DM-based IR (choi2021ilvr, ; repaint, ; kawar2021snips, ; CCDF, ; zhang2022unsupervised-pretrained, ) is proposed to exploit the intrinsic knowledge within the pre-trained diffusion models. Concretely, it is observed that pre-trained diffusion models for image generation, trained with a large number of natural images, encapsulate considerable prior knowledge about rich textures. And thus, these pre-trained diffusion models can be considered as repositories of texture information. The exploration of reusing such prior knowledge for training-free image restoration is an emerging and promising direction in the low-level vision field.
As the initial work, Choi et al.(choi2021ilvr, ) introduces the iterative latent variable refinement (i.e., ILVR) method, which leverages the unconditional diffusion model to enable the training-free conditional generation for image SR and image translation. The pivotal innovation of ILVR involves substituting the low-frequency components in denoised output with their counterpart from the reference image. This substitution process, illustrated in Fig. 4, ensures structure and semantic consistency between the generated and reference images, thereby facilitating the conditional generation. Particularly, given a reference image $y$ (e.g., the distorted image in IR), at time step $t$, the ILVR predicts the denoised result for time t1t-1 using the following formula:
@@
Here, $\Phi_{N}$ denotes the low-pass filter designed to bypass the low-frequency component from input, and $y_{t-1}\sim q(y_{t-1}|y)$ is the diffused state of $y$ at t1t-1 steps. Following ILVR, most zero-shot DM-based IR methods (dps, ; iigdm, ; DDRMDRM, ; ddnm, ; fabian2023diracdiffusion, ; feng2023score, ; copaint, ) predominantly focus on enhancing the refinement strategies within the sampling process, thereby training-free.
@@
Motivated by SR3 (SR3, ), numerous studies have endeavored to optimize the supervised DM-based IR framework, focusing on enhancing the condition strategy and exploring potential and more efficient generation spaces. With respect to the condition strategy, we categorize these studies into three types based on the conditions: 1) low-quality reference image, 2) pre-processed references, and 3) revising diffusion process. Regarding generation space, the supervised DM-based IR methods can be classified into three distinct groups: image space, residual space, and latent space. If not mentioned, the majority of studies generate the restored images within image space, where the structures and textures are required to be generated directly. In contrast, the residual-space diffusion model focuses on reconstructing the residuals between the low-quality image and its corresponding high-quality image, which simplifies the complexity of generating the whole image. Latent-based methods utilize a well-designed encoder to transform the image into a compact latent space for generation, thereby improving the generation efficiency. This section will elucidate existing studies on supervised DM-based IR in terms of the above three conditional strategies and the last two generation spaces.
@@ 3.3.1 Condition with Low-quality Reference Image
As highlighted in Sec. 3.1, incorporating distorted images as conditions is indispensable and crucial for supervised DM-based IR (chen2023image-PromptSR, ; wang2023learning-C2F-DFT, ; li2024blinddiff, ). The SR3 method has shown that substantial performance can be achieved through a simple concatenation operation. It utilizes a direct concatenation of low-quality reference images with the denoised result at the t1t-1 step as the condition for noise prediction at the $t$ step. With the same condition strategy, Saharial et al.(saharia2022palette, ) propose a unified diffusion model, termed Palette, for image-to-image translation tasks, which achieves excellent performance on image colorization, inpainting, uncropping, and JPEG artifact removal. Furthermore, they investigate the effects of different optimization objectives for sample diversity and highlight the pivotal role of self-attention within U-Net for the diffusion model. Despite their effectiveness, the aforementioned methods are constrained since they only support a fixed resolution for IR once trained. To adapt the diffusion model for the real-world IR with an arbitrary size, Özdenizci et al.(weather-diff, ) divide the degraded image and corresponding sampling results $x_{t}$ into several overlapped patches, and then utilize the patch-wise concatenation as the inputs of the diffusion model for noise prediction. Additionally, to address the inconsistency problem caused by different sampling patches in the overlapped region, this work introduces the mean estimated noise for each pixel within the overlapped region. To enhance the quality of generated images, Ho et al.(CDM, ) introduce the three-layer cascaded diffusion model based on the SR3 backbone. The first diffusion model is exploited to achieve the class-conditioned low-resolution image generation, and two additional diffusion models are cascaded to super-resolve the low-resolution generated image, resulting in a higher-resolution and more realistic generated image. In addition to the DDPM-based approaches, there is another work (sde-sr, ) that explores the various variants of the continuous diffusion model SDE (SDE, ) for the face super-resolution using predictor-corrector sampling.
@@ 3.3.2 Condition with Pre-processed References
Even though direct concatenation with the low-quality image shows promising performance, the artifacts in low-quality images inevitably cause harmful effects on the generation of the diffusion model, especially for severe and diverse distortions. To mitigate this issue, several studies (yin2023cle-CLE-Diffusion, ; wang2023lldiffusion-LLDiffuion, ; shang2023resdiff, ; niu2023cdpmsr, ; intro-densely-13, ; moser2023waving, ; guo2023boundary-BCDiff, ; wang2024promptrr, ; wu2023hsr-HSR-diff, ) strive to enhance the condition by preprocessing low-quality images with either jointly trained modules or pre-trained restoration networks. As depicted in Fig. 5, these works can be grouped into two categories based on the preprocessing strategies, i.e., condition with pre-processed reference image and feature.
@@
Pre-processed reference image. To alleviate the side effects of artifacts in low-quality images, CDPMSR (niu2023cdpmsr, ) exploits existing super-resolution models, e.g., RCAN (zhang2018image-RCAN, ), SwinIR (swinIR-transformer, ), EDSR (EDSR-9, ), to enhance the low-quality image, thereby providing a high-quality and more reliable condition for diffusion model. Additionally, it eschews the stochastic sampling during the reverse process in favor of a deterministic denoising process, yielding superior image quality and faster inference. Notably, the pre-processing reference image can serve as not only an enhanced condition but also an initially well-restored image. Therefore, ResDiff (shang2023resdiff, ) utilizes a pre-trained CNN to generate a low-frequency content-abundant image as the initially restored image and exploits the conditional diffusion model to further generate the residuals between the pre-processed distorted image and its corresponding clean image. In contrast, DehazeDDPM (yu2023high-HazeDDPM, ) tries to add a physics-aware image as an auxiliary condition to achieve complex dehazing task. Thus they generate the transmission map, and haze-free image $I$ through physical modeling. Then, these preprocessed images are then incorporated into diffusion through dynamic fusion.
@@
Pre-processed reference feature. Another popular line is to use the feature of the reference image as the condition for the diffusion model (varanka2024pfstorer, ; guo2023shadowdiffusion, ; IDM, ; jin2022des3, ; xiao2023ediffsr, ; wang2023lldiffusion-LLDiffuion, ; liu2024diff-plugin, ). IDM (IDM, ), striving for continuous image super-resolution, first extracts the initial features of the low-resolution image with EDSR (EDSR-9, ). Then, the initial features are downsampled to multiple scales, which are utilized as the conditions for different upsampling layers in the diffusion model, with the intent of refining the implicit representation. In contrast, ShadowDiffusion (guo2023shadowdiffusion, ) leverages a pre-trained transformer backbone to extract the degradation prior (i.e., the degradation-related features) from the distorted reference image. This extracted degradation prior is exploited as the auxiliary to refine the generated shadow mask and serves as the condition for shadow-free image generation.
@@
Figure 6: The architecture of IR-SDE (IR-SDE, ), where the forward diffusion process approximately models the image degradation processes. The final output of the forward process is equal to a low-quality image with random noise.
@@ 3.3.3 Condition by Revising Diffusion Process
Notably, the above-mentioned supervised DM-based IR methods introduce the conditions by modifying the network while preserving the diffusion process from DDPM (DDPM, ). However, this requires the generation process to start from the noise. To obviate this, some studies (IR-SDE, ; luo2023refusion, ; luo2023controlling-DACLIP, ; yang2023difflle-DiffLLE, ; zhao2023towards-IDM, ; wang2023sinsr, ; wang2024sam-diffsr, ; HFS-SDE, ; INDI, ; xiao2023ediffsr, ; mirza2023learning-fourier-frequency-FDB, ) condition the diffusion model by modifying the diffusion process, such that the diffused output $x_{T}$ (i.e., the start point during the reverse process) approximates the low-quality image corrupted with few Gaussian noises. As shown in Fig. 6, Luo et al.(IR-SDE, ) modify the forward process with mean-reverting SDE to model the unified degradation process of IR as:
@@
where $\theta_{t}$ and $\sigma_{t}$ are the time-dependent parameters. $\mu$ denotes the distorted image and $x$ represents its corresponding high-quality image. With the mean-reverting SDE, this work successfully models the image degradation and restoration processes with the modified forward and reverse processes, respectively. This avoids the generation from pure noise and achieves better resto-ration performance. Building on IR-SDE (IR-SDE, ), the same team introduces the Refusion (luo2023refusion, ), which further refines IR-SDE (IR-SDE, ) by optimizing aspects such as network architecture, noise level, and denoising steps, etc. To reduce the computation cost, Refusion introduces the U-Net compression strategy, thereby enabling efficient sampling in the latent space.
With the same purpose, Xie et al (xie2023diffusion, ) redefine the diffusion process such that the sampling starts from the noisy image. Taking into account the diversity of noise, they derive three separate diffusion processes for the removal of Gaussian, Gamma, and Poisson noises, respectively. In contrast, InDI (INDI, ) introduces the continuous diffusion process as:
@@
where $x$ and $y$ are the high-quality images and their corresponding low-quality counterparts. $\epsilon$ is a small constant. It can be interpreted as the step-wise interpolation of high-quality and low-quality images at time step $t$, which decomposes the original single-step prediction of supervised image restoration into several small steps, effectively circumventing the regression-to-the-mean effects often found in conventional supervised image restoration. Different from the diffusion process in the spatial domain, HFS-SDE (HFS-SDE, ) reformulate the diffusion process for magnetic resonance (MR) reconstruction at frequency space. In this approach, the forward process progressively adds noise into high-frequency space, resulting in the final $x_{T}$ that is composed of high-frequency noise and low-frequency data. During the reverse process, HFS-SDE employs Predictor-Corrector (PC) method (SDE, ) for sampling.
Figure 7: The architecture of Predict and Refine model (deblur-DPM, ). The predictor $g_{\theta}$ generates the initial prediction of a clean image, while residual information is modeled by diffusion process.